Question 12 How many sales are required to earn a target after-tax net income of $88000 if total fixed costs are $96000, the contribution margin ratio is 40%, and the tax rate is 20%? O $7040000 O $%240000 $515000 $1340000
Answer #1

Target before-tax net income=$88000/(1-tax rate)

=$88000/(1-0.2)=$110,000

Target Contribution margin=Target before tax net income+Fixed costs

=(110,000+96000)=$206000

Hence target sales=$206000/Contribution margin ratio

=$206000/0.4

=-$515,000.